### v1

# Fine-tuning an ECG Foundation Model to Predict Coronary CT Angiography Outcomes

ABSTRACT
Coronary artery disease (CAD) remains a major global health burden. Accurate identification of the culprit vessel and assessment of stenosis severity are essential for guiding individualized therapy. Although coronary CT angiography (CCTA) is the first-line non-invasive modality for CAD diagnosis, its dependence on high-end equipment, radiation exposure, and strict patient cooperation limits large-scale use. With advances in artificial intelligence (AI) and the widespread availability of electrocardiography (ECG), AI-ECG offers a promising alternative for CAD screening. In this study, we developed an interpretable AI-ECG model to predict severe or complete stenosis of the four major coronary arteries on CCTA. On the internal validation set, the model’s AUCs for the right coronary artery (RCA), left main coronary artery (LM), left anterior descending artery (LAD), and left circumflex artery (LCX) were 0.794, 0.818, 0.744, and 0.755, respectively; on the external validation set, the AUCs reached 0.749, 0.971, 0.667, and 0.727, respectively. Performance remained stable in a clinically normal-ECG subset, indicating robustness beyond overt ECG abnormalities. Subgroup analyses across demographic and acquisition-time strata further confirmed model stability. Risk stratification based on vessel-specific incidence thresholds showed consistent separation on calibration and cumulative event curves. Interpretability analyses revealed distinct waveform differences between high- and low-risk groups, highlighting key electrophysiological regions contributing to model decisions and offering new insights into the ECG correlates of coronary stenosis.
KEYWORDS
Electrocardiogram, Coronary CT Angiography, Artificial Intelligence, Deep Learning, Coronary Heart Disease
INTRODUCTION
Coronary atherosclerotic heart disease (CAD) is one of the leading causes of death and disability worldwide, placing a heavy burden on public health systems and socioeconomic systems 1, 2, 3. Accurately identifying the responsible vessel in CAD, assessing lesion severity, and guiding individualized treatment decisions remain core challenges in clinical practice 4.
Currently, coronary CT angiography (CCTA) is recommended as the first-line non-invasive modality for diagnosing CAD, providing a detailed assessment of coronary anatomy and stenosis severity, while serving as a crucial tool for pre-procedural planning and strategy guidance for percutaneous coronary intervention (PCI)5, 6. However, although CCTA is a non-invasive examination, the procedure requires sophisticated equipment and image post-processing techniques, and is highly dependent on patient heart rate control and respiratory coordination7. Furthermore, CCTA carries certain radiation exposure risks8 and may cause contrast-induced kidney injury9, limiting its application, especially in patients with renal insufficiency. Therefore, exploring a rapid, low-cost, non-invasive alternative that accurately reflects coronary artery lesions is of significant clinical importance for optimizing treatment processes, improving the efficiency of medical resource utilization, and enhancing patient outcomes.
Electrocardiography (ECG), as a non-invasive, convenient, and real-time detection method, has unique advantages in the early identification and risk stratification of cardiovascular diseases. Changes in ECG waveforms, such as ST segment shift and T wave inversion, have been widely used in the preliminary diagnosis of acute coronary syndrome (ACS)10, 11, 12. Meanwhile, with the rapid development of artificial intelligence (AI), especially deep learning technology, AI is being widely used in clinical research 13, 14, 15, 16, 17, 18. Utilizing AI to perform deep feature mining of ECG signals to predict the pathogenesis of cardiovascular diseases has become a research hotspot.19, 20, 21, 22, 23. However, existing AI-ECG studies on CAD-related coronary artery disease using CCTA mostly focus on whether there is screening for obstructive coronary artery disease. For example, Choi et al.’s study mainly targeted the detection of obstructive coronary artery disease24, while Park et al.’s study focused on predicting obstructive lesions in patients with stable angina25. Although some studies have attempted to use ECG and AI models to locate myocardial infarction (MI) sites, their research subjects are mostly scar tissue from previous myocardial infarctions, lacking systematic research on the localization of the responsible vessel during acute coronary syndrome and the identification of coronary artery occlusion on normal ECG 26.
To address the aforementioned shortcomings, this study integrated clinical data from multiple centers, including Peking University People’s Hospital and the Second Hospital of Tianjin Medical University. Employing a transfer learning strategy, we fine-tuned ECGFounder27, 28, the world’s largest ECG foundation model trained for predicting future new-onset MI. We constructed a predictive model that can directly infer the diagnostic results of severe stenosis and complete occlusion shown by CCTA from ECG signals, and further implemented a stratified assessment of the risk of future new-onset stenosis in major coronary arteries. Simultaneously, we systematically evaluated the model’s performance in a population with normal ECG, providing a new technical approach for the early screening and risk identification of occult CAD. Furthermore, this study deeply analyzed the changing characteristics of ECG waveforms in different risk populations, revealing potential signal patterns related to coronary artery stenosis from an electrophysiological perspective, providing a new perspective for understanding its pathogenesis, and further expanding the application boundaries of ECG in the precise diagnosis of coronary artery stenosis.
Overall, the principal innovations and contributions of this study are summarized as follows.
  1. 1.
    We developed an interpretable AI-ECG model capable of directly predicting CCTA-defined severe coronary stenosis and complete occlusion from ECG signals. The predictive performance and probability calibration of the model were systematically evaluated using receiver operating characteristic (ROC) curves, box plots of predicted probabilities across different stenosis severities with statistical testing, and calibration curves.
  2. 2.
    We further investigated the model’s ability to identify CCTA-defined severe stenosis and complete occlusion in individuals with apparently normal ECGs, demonstrating its potential value for the detection of concealed CAD.
  3. 3.
    By integrating model-predicted probabilities with observed incidence rates, we stratified individuals into high-risk and low-risk groups and constructed cumulative incidence curves, enabling dynamic risk stratification for future coronary events.
  4. 4.
    Through explainable analysis, we comprehensively compared ECG waveform characteristics between model-defined high-risk and low-risk populations, providing novel insights into the electrocardiographic manifestations associated with coronary artery stenosis and offering a new perspective for clinical interpretation.
RESULTS
Baseline Character of Dataset
This study used 2,595 CCTA examinations and 4,620 paired ECG data from 2,323 patients at Peking University People’s Hospital, and 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Hospital of Tianjin Medical University. We used the data from Peking University People’s Hospital as the internal dataset and the data from the Second Hospital of Tianjin Medical University as the external dataset to validate the model’s generalization ability on multicenter data.
Table 1 shows the baseline characteristics of the internal and external data used in the study.
Table 1: Baseline Characteristics of Internal and External Cohorts
Characteristic Peking University People’s Hospital (Internal) The Second Hospital of Tianjin Medical University (External)
No. of Patients 2323 135
Age (years)
      Mean $\pm$ SD $63.7\pm 10.1$ $62.6\pm 14.2$
      Median (Range) 64.0(21.092.0)64.0\ (21.0-92.0) 65.0(28.093.0)65.0\ (28.0-93.0)
Sex (n (%))
      Male $1042\ (55.0\%)$ $80\ (61.1\%)$
      Female $853\ (45.0\%)$ $51\ (38.9\%)$
Development and Multi-Center Validation of AI-ECG Model
We used patient ID as the grouping criterion and employed a five-fold stratified grouping cross-validation strategy to train and validate the AI-ECG model while ensuring a balanced distribution of multi-labels across each fold. To systematically evaluate the model’s generalization performance on a multi-center dataset and its predictive ability in patients with clinically diagnosed normal ECGs, this study selected the best-performing fold result from the internal validation set, the overall result from the external validation set, and the prediction result of the best-performing fold for patients with ECG diagnoses of normal. Three ROC curves were plotted, as shown in Figure 1.
The results show that, in the internal validation set, the predicted AUCs for the right coronary artery (RCA), left main coronary artery (LM), left anterior descending artery (LAD), and left circumflex artery (LCX) are 0.794, 0.818, 0.744, and 0.755, respectively; while in the external validation set, the corresponding AUCs are 0.749, 0.971, 0.667, and 0.727. These results demonstrate that our AI-ECG model exhibits stable and superior discriminative performance in identifying severe stenosis and complete occlusion of the four major coronary arteries. The external validation results are generally consistent with the internal validation results, further confirming the model’s good generalization ability across different central datasets.
Furthermore, we evaluated the model’s predictive ability in the internal validation set for individuals with clinically normal ECGs. The results showed that in this subgroup, the predictive AUCs for RCA, LM, LAD, and LCX were 0.793, 0.896, 0.707, and 0.765, respectively. These results are largely consistent with the model’s performance in the full population within the internal validation set, indicating that even in individuals with superficially normal ECGs, the AI-ECG model can effectively identify the potential risk of severe coronary artery stenosis, highlighting its potential application value in screening for occult CAD.
Meanwhile, we calculated the 95% confidence interval for the AUC index of each blood vessel. The range of this interval effectively reflects the statistical reliability and stability of the model’s predictive performance, further supporting the credibility of the model’s discriminative ability.
(a) Internal Validation
(b) External Validation
(c) Normal ECG Prediction
Figure 1: ROC curves for model performance evaluation. (a) ROC curve validated using internal data from Peking University People’s Hospital. (b) External validation ROC curve of the Second Hospital of Tianjin Medical University. (c) Predicted ROC curves of normal electrocardiograms from internal data.
The study further conducted subgroup analyses in the internal validation set to assess the model’s stability across different population characteristics and time conditions. After stratifying the population according to age (<65years<65years vs. $\geq 65years$), sex, and the time interval between ECG and CCTA examinations ($\leq 3hours$ vs. >3hours>3hours), the model’s discriminative performance remained consistent across all subgroups.
As shown in Figure 2, the AUC levels of the model remained generally stable across subgroups, with no significant performance decline.
Specifically, the model exhibited a relatively higher AUC in the subgroup where the interval between ECG and CCTA examinations was greater than 3 hours; it also demonstrated superior discriminative performance in the younger group (under 65 years of age). Between the male and female subgroups, the model’s predictive performance was largely consistent, with no significant differences.
These results further demonstrate that the model possesses good stability and robustness under different population characteristics and testing conditions, further supporting its potential application value in diverse clinical scenarios.
Figure 2: Subgroup analyses. Subgroup analysis results demonstrate that the model’s predictive performance remains robust across different age groups, genders, and time intervals.
To further validate the correlation between the model’s predicted probabilities and the actual degree of stenosis, we plotted a box plot as shown in Figure 3. The box plot illustrates the predicted probability distribution of the four major coronary arteries at four degrees of stenosis (normal, mild, moderate, and severe). All vessels exhibit a consistent monotonically increasing trend, indicating that the model assigns a higher risk probability as the degree of stenosis increases.
Figure 3: Distribution of model-predicted probabilities across different degrees of coronary stenosis. Box plots illustrate the predicted probability distributions for the four major coronary arteries (LAD, LCX, LM, and RCA) across four stenosis categories (normal, mild, moderate, and severe).
Spearman rank correlation analysis confirmed a statistically significant positive correlation between the predicted probability and the degree of stenosis in all coronary arteries ($p<0.001$), with the left anterior descending artery (LAD) showing the strongest correlation ($\rho=0.435$). Notably, the absolute predicted probability of the left main coronary artery (LM) (range: 0.0–0.2) was significantly lower than that of other vessels (range: 0.0–0.8), reflecting the low prevalence of LM lesions in the training population. However, the model still maintained significant discriminative power for LM ($\rho=0.307$).
Model’s Decision-making Ability in Coronary Artery Stenosis Risk Stratification
For the four major coronary arteries, this study established differentiated risk thresholds based on the epidemiological characteristics of severe stenosis and complete occlusion in each vessel, and divided patients into high-risk and low-risk groups based on model-predicted probabilities. Specifically, for LAD, LCX, and RCA, which have relatively high incidence rates, a predicted probability $\geq 15\%$ was defined as the high-risk threshold, referencing the pre-test probability stratification model proposed in the 2019 ESC guidelines. For LM, which has a significantly lower incidence rate, a predicted probability $\geq 1\%$ was set as the high-risk criterion based on an incidence rate enrichment strategy, in order to improve the sensitivity of identifying rare but clinically significant high-risk lesions.
To verify the reliability of the model’s predictions, we plotted the model’s prediction calibration curves on an internal dataset, as shown in Figure 4.
Figure 4: Calibration curve of model prediction results. The calibration curves demonstrate the model’s reliability in predicting each vessel, and all four coronary arteries exhibited low Brier scores.
The calibration curves show that the model demonstrates good reliability in predicting all four coronary arteries. The Bridge scores for RCA, LM, LAD, and LCX are 0.112, 0.030, 0.166, and 0.106, respectively. Notably, for severe and complete occlusion of the LM, the model did not output a very high prediction probability due to the extremely low incidence of this lesion in the general population (approximately 0.1%), which is consistent with epidemiological distribution.
To clearly demonstrate its performance, we magnified the LM subplot locally (between 0 and 0.3). The results show that although the predicted probabilities are concentrated in the low range, the high-risk group still exhibits a clear upward trend in risk, and the Brier Score (0.030) is extremely low, proving that the model has extremely high accuracy in excluding the vast majority of healthy people (High Negative Predictive Value), and does not produce a large number of false positives due to blindly pursuing high probabilities.
To further validate the model’s decision-making ability in clinical risk stratification, we plotted the cumulative event occurrence curves based on the model’s predicted risk stratification, as shown in Figure 5. Using the ECG monitoring time as the follow-up starting point, we defined the occurrence of severe stenosis and complete occlusion as the event endpoints, with a follow-up period of one year. Individuals who did not develop severe stenosis or complete occlusion during the one-year follow-up period were defined as having no events. By comparing the cumulative event occurrence curves of high-risk and low-risk populations, we assessed the model’s ability to distinguish future coronary events, thereby validating its clinical risk stratification value over time.
Figure 5: Cumulative event occurrence curve based on model-predicted risk stratification. The cumulative event occurrence curve shows that the stratification results for each coronary artery exhibit a good separation trend and significant statistical differences.
As shown in Figure 5, the cumulative event occurrence curves of the high-risk and low-risk groups in the four major coronary arteries exhibit a clear and stable separation trend, demonstrating the model’s good ability to distinguish between different risk levels. The differences between the two groups were statistically significant (log-rank test, all P < 0.001), further validating the model’s discriminative power in coronary event risk stratification and its potential clinical application value.
Interpretability Analysis Results
To further reveal the key ECG regions that the model focuses on in risk assessment and stratification, we conducted an interpretability analysis based on the waveform level. After dividing the population into high-risk and low-risk groups according to the model’s predicted probabilities, we extracted normalized individual heartbeat segments from each patient’s raw ECG signals and aligned and standardized them. Subsequently, within each risk group, we calculated the mean morphology and standard deviation of the heartbeat waveforms in each lead and presented them in a multi-lead visualization.
As shown in Figure 6, by comparing the differences in waveform morphology between high-risk and low-risk groups, the electrocardiographic feature patterns on which the model distinguishes different coronary artery risk states are intuitively revealed, providing interpretable evidence for understanding its discrimination mechanism and potential electrophysiological basis.
Figure 6: Interpretability analysis results. The results reveal the key ECG regions that the model focuses on during prediction and stratification.
DISCUSSION
This study constructs an interpretable AI-ECG model for predicting severe stenosis and total occlusion outcomes in CCTA using ECG. The model is based on ECGFounder, which has been trained for predicting new-onset MI, and fine-tuned using transfer learning to address shortcomings in existing studies regarding the study population and predicted outcomes. The model’s predictive performance was systematically validated using ROC curves, subgroup analysis, and box plots. Furthermore, based on the incidence distribution of major coronary artery lesions, a prediction probability threshold was set to define high-risk groups, and the model’s risk stratification capability was further evaluated using calibration curves and cumulative event occurrence curves.
The main innovations of this study are reflected in the following aspects: First, the proposed model can directly extract clinically relevant information on severe and complete CCTA blockage from the raw ECG signal, and still shows good consistency in ECGs that are clinically diagnosed as normal, suggesting the potential application value of the model in screening for occult CAD; Second, the study systematically evaluated the model’s performance in risk stratification, and the results show that it has the ability to assist in clinical decision stratification; Finally, through comparative analysis of ECG waveform characteristics between high-risk and low-risk populations, this study attempts to explore potential pathological differences at the signal level, providing a new perspective for the study of CAD mechanisms.
While this study demonstrates some innovation in model construction and clinical application exploration, several limitations remain. First, the data primarily came from two centers, and the sample size and population diversity need further expansion, potentially limiting the model’s generalization ability. Second, due to the low incidence of severe and complete LM obstruction, the AUC of LM was skewed in the external validation set and among individuals with normal ECG during model training and validation, suggesting that the results may be affected by sample imbalance and require careful interpretation.
In future research, we will incorporate large-scale data from more centers to further improve the model’s stability and generalization ability. Simultaneously, we will conduct more refined analyses of LM lesions, such as combining spatial localization information from anterior and posterior leads to explore ECG characteristics corresponding to different anatomical regions, thereby improving the model’s ability to identify specific coronary artery lesions.
METHODS
This study aimed to construct an interpretable AI-ECG model for predicting CCTA results from ECG, thus providing a potential alternative for non-invasive risk assessment of CAD. The model was developed based on 2,595 CCTA examinations and 4,620 paired ECG data from 2,323 patients at Peking University People’s Hospital, China. External validation was performed on 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Hospital of Tianjin Medical University, China, to systematically evaluate the model’s generalization performance across different centers and populations.
To assess the clinical applicability of the model, this study constructed a systematic performance evaluation framework, comprehensively validating the model from multiple dimensions, including discriminative ability, probabilistic reliability, and clinical application potential. First, ROC curves were plotted to evaluate the model’s discriminative performance in identifying severe and complete occlusion in CCTA. Box plots of the model’s predicted probabilities for different degrees of stenosis were then used to quantitatively analyze the model’s ability to distinguish the severity of coronary artery stenosis. Furthermore, based on the model’s predicted probabilities and the actual incidence of severe coronary artery occlusion, the study population was stratified into high-risk and low-risk groups. Simultaneously, calibration curves were used to systematically test the reliability of the model’s predicted probabilities, and cumulative event occurrence curves for different risk groups were plotted to evaluate the model’s stratified predictive ability for future coronary events.
Furthermore, this study systematically evaluated the model’s classification performance across different age groups, sexes, and subgroups with varying ECG and CCTA examination time differences. We further explored its ability to identify severe coronary artery stenosis on normal ECGs, thereby validating the model’s potential application value in screening for occult lesions in coronary artery disease (CAD). This evaluation process extends from classification performance to probabilistic reliability, and further to clinical decision support potential, constructing a complete and coherent clinical performance evaluation pathway for the model.
To enhance the interpretability of the model, this study further compared the differences in single heartbeat waveforms in the twelve leads among different risk groups classified by the model, analyzed the key waveform regions that the model focuses on and their association with coronary artery lesions, thereby exploring the model’s discrimination criteria from an electrophysiological perspective, providing intuitive evidence for understanding the mechanism of AI-ECG in CAD identification, and providing a new research perspective for related clinical studies.
Data Sources and Labeling Strategy
This study included 2,595 CCTA examinations and 4,620 paired electrocardiogram (ECG) data from 2,323 patients at Peking University People’s Hospital in China for model development, and external validation was performed on 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Affiliated Hospital of Tianjin Medical University in China. To avoid bias caused by potential ECG bias due to clinical interventions such as catheterization lab procedures following CCTA examinations, the ECG examination time was strictly limited to be earlier than the corresponding CCTA examination time during data cleaning. This study has been approved by the Ethics Committee of Peking University (Approval No.: IRB00001052-23071), and the Ethics Committee has approved the waiver of the informed consent form.
Based on CCTA examination reports, under the guidance of professional clinicians, the original diagnostic text was structured, cleaned, and standardized using a large language model to extract stenosis information for the RCA, LAD, LCX, and LM. Based on clinical consensus, "normal, mild stenosis, and moderate stenosis" were categorized into the non-severe stenosis group, while "severe stenosis and complete occlusion" were categorized into the severe lesion group. This stratification method aligns with clinical criteria for determining lesions with significant prognostic value.
Meanwhile, the ECG diagnostic texts were uniformly cleaned and recoded, and the study population was further divided into two categories: "normal ECG" and "abnormal ECG" to support subsequent stratified analysis of the ability to identify occult coronary artery lesions.
Model Architecture and Training Strategy
This study fine-tuned the ECGFounder model trained for future new-onset MI using a transfer learning strategy 27, 28. An AI-ECG model based on a multi-task deep learning framework was constructed to jointly predict severe stenosis and complete occlusion of four major coronary arteries (RCA, LM, LAD, LCX). The model uses a Net1D network structure with 12-lead ECG signals as input.
To address the gradient conflict and weight imbalance issues between different tasks during multi-task learning, an adaptive weighting strategy based on task uncertainty was introduced 29. Specifically, for each task $t$, the loss function was weighted by a learnable uncertainty parameter $\sigma_{t}$, and the overall objective was defined as:
$\mathcal{L}=\sum_{t=1}^{T}\left(\frac{1}{2\sigma_{t}^{2}}\mathcal{L}_{t}+\log\sigma_{t}\right),$ (1)
where $\mathcal{L}_{t}$ denotes the task-specific loss and $\sigma_{t}$ represents the task uncertainty, which is optimized jointly with model parameters. This formulation enables dynamic balancing of multiple tasks during training.
In addition, the Projected Conflicting Gradient (PCGrad) algorithm was applied to further mitigate gradient interference between tasks 30. For two task gradients $g_{i}$ and $g_{j}$, when their dot product is negative, indicating conflicting optimization directions, the gradient was adjusted as:
$g_{i}\leftarrow g_{i}-\frac{g_{i}^{\top}g_{j}}{\|g_{j}\|^{2}}g_{j},\quad\text{if }g_{i}^{\top}g_{j}<0,$ (2)
thereby projecting conflicting gradients onto a non-conflicting direction to improve convergence stability.
The optimization process employed the AdamW optimizer combined with warmup and cosine decay strategies to dynamically adjust the learning rate, enhancing training robustness and generalization performance.
Data partitioning employed a five-fold stratified grouped cross-validation method, using patient ID as the grouping criterion to ensure that data from the same patient did not appear simultaneously in both the training and validation sets, while maintaining the balance of multi-label distribution. The model used the validation set Macro-AUC as the primary performance monitoring metric, and the weights of the best-performing model in each fold were stored.
After reading the 12-lead ECG signal, lead-by-lead z-score standardization was first performed to eliminate the impact of inter-individual amplitude differences on model training. To enhance the model’s robustness to signal perturbations encountered in real-world clinical scenarios, an online data augmentation strategy was employed during training to generate augmented ECG samples via controlled random perturbations.
Specifically, each ECG signal was subjected to stochastic transformations, including random temporal shifting, amplitude scaling, additive Gaussian noise injection, and local random occlusion. Temporal shifting was implemented by rolling the signal along the time axis within a range of ±10% of the signal length, simulating beat misalignment. Amplitude scaling was performed by multiplying the signal with a random factor sampled from a uniform distribution in the range [0.9, 1.1], mimicking inter-beat voltage variation. Gaussian noise with zero mean and variance proportional to the augmentation intensity was added to emulate physiological and acquisition noise, while local random occlusion was applied by setting a contiguous segment (2%–10% of signal length) to zero, simulating signal dropout or artifact interference.
The augmentation strength was modulated by an epoch-dependent cosine annealing function:
$\alpha(e)=0.5+0.5\cos\left(\frac{\pi e}{E}\right),$ (3)
where $e$ denotes the current training epoch and $E$ represents the total number of training epochs. This formulation ensured that augmentation intensity gradually decreased from 1.0 to 0.5 throughout training, enabling the model to initially learn robust invariant representations under strong perturbations and subsequently focus on stable morphological features in later stages, thereby improving generalization while avoiding excessive distortion of physiologically relevant ECG patterns.
During training, both the original ECG signal and its augmented counterpart were simultaneously fed into the network, encouraging the model to learn consistent representations across perturbed signal variations.
The training strategy described above, while ensuring data isolation at the patient level, enables the model to make stable joint predictions of multivascular lesions and further outputs continuous probability results, providing a methodological basis for subsequent risk stratification analysis, calibration assessment and clinical decision support.
Evaluation Methods
To systematically evaluate the predictive performance of the model, the ROC curves were plotted in both internal and external validation sets, and the 95% confidence intervals for each AUC were calculated to quantify the classification and discrimination capabilities of the model.
Based on different true diagnostic categories in CCTA, box plots of the model’s predicted probabilities for corresponding patients were generated, and the Spearman rank correlation test was used to assess the trend relationship between different coronary artery stenosis grades and the model’s predicted probabilities. By encoding the stenosis degree into ordered numerical codes ("no obvious stenosis, mild stenosis, moderate stenosis, severe stenosis"), the monotonic correlation between these codes and the model’s output probability was analyzed to determine whether the model’s predicted probability showed a significant upward trend with increasing stenosis degree, thus validating the model’s discriminative power and reliability from a probability distribution perspective.
Furthermore, we conducted subgroup analyses on the internal validation set, stratifying the population based on age (<65<65 years vs. $\geq 65$ years), sex, and the time interval between ECG and CCTA examinations ($\leq 3$ hours vs. >3>3 hours). We systematically compared the differences in model performance among different subgroups to assess the model’s robustness under various population characteristics and time conditions. Finally, for patients with normal ECG diagnoses, we further analyzed the model’s prediction results to validate the potential application value in detecting occult CAD.
Clinical Risk Stratification
Based on the differences in the incidence of the four major coronary arteries (LAD, LCX, LM, RCA) in the real-world population, we divided the vessels into two groups: relatively high incidence and extremely low incidence, to set differentiated model prediction thresholds. Referring to the results of Bergström et al.’s study based on the large-scale SCAPIS cohort ($n=25,182$), the anatomical distribution of significant coronary artery stenosis differs greatly in the general population without known coronary artery disease31. Specifically, the probability of significant stenosis in the LM is extremely low, only 0.1%; in contrast, the incidence in other major vessels is approximately 20 times higher (1.9%) than in the LM. Based on this significant epidemiological difference, we grouped LAD, LCX, and RCA into the relatively high incidence group, while classifying the LM separately into the extremely low incidence group.
For blood vessels with high incidence (LAD, LCX, RCA), referring to the pre-test probability (PTP) model based on a pooled analysis of contemporary large-scale patient cohorts in the 2019 ESC guidelines, a predicted probability of $\geq 15\%$ from the AI-ECG model was defined as high risk32. For LM with extremely low incidence, considering its extremely low baseline incidence in the general population (0.1%), we adopted an incidence enrichment strategy, defining a predicted risk of 10 times the baseline incidence (i.e., $\geq 1.0\%$) as high risk33. Although this threshold is low in absolute value, it represents a highly significant relative risk ($RR=10$), aiming to significantly reduce the number of people requiring screening (NNS) in clinical screening, thereby improving diagnostic benefits.
Interpretability Analysis
The study stratifies patients according to predicted probabilities in the model and further compares the characteristic differences in the ECG waveforms between different risk groups to improve the interpretability of the AI-ECG model and reveal the electrophysiological basis of its risk assessment.
Specifically, based on the predicted probabilities corresponding to each coronary arteries and preset risk thresholds, the population is divided into high-risk and low-risk groups. R waves are then automatically detected from the raw ECG signals of each patient, and individual heartbeats are time-aligned to extract standardized heartbeat segments, which are then normalized by z-score for each lead. The heartbeat waveforms within different risk groups are summarized and calculated to obtain the average waveform and standard deviation for each lead, and visualized using a multi-lead grid.
By comparing the morphological differences in key electrophysiological features such as P waves, QRS complexes, and ST-T segments between high-risk and low-risk individuals, the waveform patterns relied upon by the model to distinguish different coronary artery risk states are intuitively revealed, thus providing intuitive evidence for understanding the model’s decision-making mechanism and its underlying physiological basis.
ACKNOWLEDGMENTS
Shenda Hong is supported by the National Natural Science Foundation of China (62102008, 62172018), CCF-Tencent Rhino-Bird Open Research Fund (CCF-Tencent RAGR20250108), CCF-Zhipu Large Model Innovation Fund (CCF-Zhipu202414), PKU-OPPO Fund (BO202301, BO202503), Research Project of Peking University in the State Key Laboratory of Vascular Homeostasis and Remodeling (2025-SKLVHR-YCTS-02).
DECLARATION OF INTERESTS
The authors declare no competing interests.